OpenStudy (aonz):

cos3x = cos2x Find general solutions and 0

OpenStudy (dan815):

oh lol why am i telling you their period

OpenStudy (dan815):

just solve for x

OpenStudy (aonz):

how??

OpenStudy (aonz):

3x = 360 + 2x or 3x = 360 - 2x

OpenStudy (dan815):